# Large Angular Momentum

### v1

@@ Acknowledgments
The work by KK is supported by the INFN special-initiative project grant GAST (Gauge and String Theories). We are grateful to Hans-Thomas Elze, Jarah Evslin, Riccardo Guida, Pietro Menotti and Arkady Vainshtein for useful comments and precious advices.
@@ Appendix A Newton’s equation for a macroscopic body
The conditions needed for the CM of an isolated macroscopic body at finite body temperatures to obey Newton’s equations have been discussed in great care by one of the present authors [1]. They are
@@
Under these conditions, the CM of an isolated macroscopic body has a unique trajectory. Newton’s equations for it follow from the Ehrenfest theorem. See Ref. [1] for various subtleties and for the explicit derivation of Newton’s equation under external gravitational forces, under weak, static, smoothly varying external electromagnetic fields, and under a harmonic-oscillator potential. Somewhat unexpectedly, the environment-induced decoherence [5, 6, 7] which is extremely effective in rendering macroscopic states in a finite-temperature environment a mixture, has been found not to be the most essential elements for the derivation of classical mechanics from quantum mechanics [1, 2].

### v3

@@ Acknowledgments
The work by K.K. is supported by the INFN special-initiative project grant GAST (Gauge and String Theories). We are grateful to Hans Thomas Elze, Jarah Evslin, Riccardo Guida, Pietro Menotti and Arkady Vainshtein for useful comments and precious advices.
@@ Appendix A Newton’s equation for a macroscopic body
The conditions needed for the CM of an isolated macroscopic body at finite body temperatures to obey Newton’s equations have been investigated in great care by one of the present authors [22]. They are
@@
Under these conditions, the CM of an isolated macroscopic body has a unique trajectory. Newton’s equations for it follow from the Ehrenfest theorem. See Ref. [22] for various subtleties and for the explicit derivation of Newton’s equation under external gravitational forces, under weak, static, smoothly varying external electromagnetic fields, and under a harmonic-oscillator potential. Somewhat unexpectedly, the environment-induced decoherence [31, 32, 33] which is extremely effective in rendering macroscopic states in a finite-temperature environment a mixture, is found not to be the most essential element for the derivation of classical mechanics from quantum mechanics [22, 23].
Appendix B A subtle face of the Stern-Gerlach experiment
The Hamiltonian is given by
$H=\frac{{\bf p}^{2}}{2m}+V\;,\qquad V=-{\boldsymbol{\mu}\cdot{\bf B}}\;,$ (B.1)
${\boldsymbol{\mu}}=\mu_{B}\,g\,{\mathbf{s}}\;,\qquad\partial B_{z}/\partial z\neq 0\;,$ (B.2)
where $\mu_{B}=\frac{e\hbar}{2m_{e}c}$ is the Bohr magneton. We recall the well-known fact that the gyromagnetic ratio $g\simeq 2$ of the electron and the spin magnitude $1/2$ approximately cancel, so $\mu_{B}$ is the magnetic moment  [^9] For illustration purpose, we use this value below, but the discussion is obviously valid for general atoms., in the case of the atoms such as $A_{g}$, where a single outmost electron provides the total spin $\tfrac{1}{2}$.
An example of the inhomogeneous field ${\bf B}$ appropriate for the Stern-Gerlach experiment is [29, 28]
${\bf B}=(0,B_{y},B_{z}),\qquad B_{y}=-b_{0}\,y,\quad B_{z}=B_{0}+b_{0}\,z\;$ (B.3)
which satisfy $\nabla\cdot{\bf B}=\nabla\times{\bf B}=0$. The constant field $B_{0}$ in the $z$ direction must be large,
$|B_{0}|\gg|b_{0}\,y|\;.$ (B.4)
in the relevant region of $(y,z)$ of the experiment. The wave function of the spin $\frac{1}{2}$ particle entering the SG magnet has the form,
$\Psi={\tilde{\psi}}_{1}({\bf r},t)|\!\uparrow\rangle+\,{\tilde{\psi}}_{2}({\bf r},t)|\!\downarrow\rangle\;.$ (B.5)
obeying the Schrödinger equation,
$i\hbar\frac{d}{dt}\Psi=H\,\Psi\;.$ (B.6)
By redefining the wave functions for the upper and down spin components as
${\tilde{\psi}}_{1}({\bf r},t)=e^{i\mu_{B}B_{0}t/\hbar}\psi_{1}({\bf r},t)\;,\qquad{\tilde{\psi}}_{2}({\bf r},t)=e^{-i\mu_{B}B_{0}t/\hbar}\psi_{2}({\bf r},t)\;,$ (B.7)
one finds that the up- and down- spin components $\psi_{1}$ and $\psi_{2}$ satisfy the separate Schrödinger equations [29]
$i\hbar\frac{\partial}{\partial t}\psi_{1}=\left(\frac{{\bf p}^{2}}{2m}-\mu_{B}b_{0}z\right)\psi_{1}\;,\qquad i\hbar\frac{\partial}{\partial t}\psi_{2}=\left(\frac{{\bf p}^{2}}{2m}+\mu_{B}b_{0}z\right)\psi_{2}\;.$ (B.8)
This is because the term $\propto-\mu_{y}B_{y}=-(g\mu_{B}b)s_{y}$ in the Hamiltonian (B.1) mixing the two components $\psi_{1,2}$ has acquired a rapidly oscillating phase factor,
±$\pm i\mu_{B}\,b_{0}y\,e^{\mp 2i\mu_{B}B_{0}t/\hbar}\;,$ (B.9)
hence can be safely neglected. The condition (B.4) is crucial here.
As explained in [28], this can be classically understood as the spin precession effect around the large constant magnetic field $B_{0}\hat{z}$, thanks to which the forces on the particle in the transverse ($\hat{x},\hat{y}$) directions average out to zero 101010 With a magnetic field $B_{0}$ of the order of $10^{3}$ Gauss, the precession frequency is of the order of $10^{11}$ in the case of the SG experiment with silver atom [28]. With the average velocity of $A_{g}$ atoms of the order of $100\,m/s$ and the size of the region of the magnetic field of the order of a few $cm$ [27], the timescale of the precession is orders of magnitude ($\sim 10^{-5}$) shorter than the time the atoms spend in the region.. The only significant force it receives is due to the inhomogeneity in $B_{z}$, (B.2), which deflects the atom in the ±$\pm\hat{z}$ direction.
From (B.8) and their complex conjugates, one can derive the Ehrenfest theorem for $\psi_{1}$ and $\psi_{2}$ separately,
$\frac{d}{dt}\langle{\bf r}\rangle_{1}=\langle{\bf p}/{m}\rangle_{1}\;,\quad\frac{d}{dt}\langle{\bf p}\rangle_{1}=-\langle\nabla(\mu_{B}B_{z})\rangle_{1}\;;$ (B.10)
$\frac{d}{dt}\langle{\bf r}\rangle_{2}=\langle{\bf p}/{m}\rangle_{2}\;,\quad\frac{d}{dt}\langle{\bf p}\rangle_{2}=+\langle\nabla(\mu_{B}B_{z})\rangle_{2}\;,$ (B.11)
where $\langle{\bf r}\rangle_{1}\equiv\langle\psi_{1}|{\bf r}|\psi_{1}\rangle$, etc. Namely, for a sufficiently compact initial wave packet, $\psi({\bf r},t)$, the expectation values of ${\bf r}$ and ${\bf p}$ in the up and down components $\psi({\bf r},t)_{1,2}$ trace respectively the classical trajectories of a spin-up or spin-down particle.
Still, even though the two subpackets might get well separated by a macroscopic distance, $\Psi$ remains in a coherent superposition of the upper and lower spin components $\psi_{1,2}$. Its pure-state nature can be verified, in a set-up known as the quantum eraser, by reconverging them by using a second magnetic field of opposite gradients, and studying their interferences [35].
A variational solution of (B.8) for Gaussian split wave packets can be found in [23].